OpenStudy (anonymous):

If x =3 and y = -2 , then y4 - 4x = ?

OpenStudy (anonymous):

-20

OpenStudy (anonymous):

(-2)^4 - 4(3) = 16-12 = 4 -->ans:)

OpenStudy (anonymous):

is it y^4 or y4 (same as 4y)?

OpenStudy (anonymous):

based on wat is written, i m guessing it is y^4...

OpenStudy (anonymous):

then a medal for you

OpenStudy (anonymous):

haha thx:) but if it was supposed to be 4y, u would be right... anyway, it is up to the asker to decide which is the answer so, medal 4 u too:)